Consider the population of the values (1,2,3,4).




A. List all the possible samples of size 2 with replacement




B. Compute the mean of each sample.




C. Identify the probability of each sample.




D. Compute the mean of the sampling distribution of the means.

1
Expert's answer
2022-05-12T02:30:07-0400

A.We have population values 1,2,3,4, population size N=5 and sample size n=2.

The number of possible samples which can be drawn with replacement is

"N^n=4^2=16."

"S=\{(1,1), (1,2), (1,3), (1,4),"

"(2,1), (2,2), (2,3), (2,4),"

"(3,1), (3,2), (3,3), (3,4),"

"(4,1), (4,2), (4,3), (4,4)\}"

B.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 1,1 & 2/2 \\
 \hdashline
 2 & 1,2 & 3/2 \\
 \hdashline
 3 & 1,3 & 4/2 \\
 \hdashline
 4 & 1,4 & 5/2 \\
 \hdashline
 5 & 2,1 & 3/2 \\
 \hdashline
 6 & 2,2 & 4/2 \\
 \hdashline
 7 & 2,3 & 5/2 \\
 \hdashline
 8 & 2,4 & 6/2 \\
 \hdashline
 9 & 3,1 & 4/2 \\
 \hdashline
 10 & 3,2 & 5/2 \\
 \hdashline
 11 & 3,3 & 6/2 \\
 \hdashline 
 12 & 3,4 & 7/2 \\
 \hdashline 
 13 & 4,1 & 5/2 \\
 \hdashline 
 14 & 4,2 & 6/2 \\
 \hdashline 
 15 & 4,3 & 7/2 \\
 \hdashline 
 16 & 4,4 & 8/2 \\
 \hdashline 
\end{array}"



C.

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 2/2 & 1/16 & 2/32 & 4/64 \\
 \hdashline
 3/2 & 2/16 & 6/32 & 18/64 \\
 \hdashline
 4/2 & 3/16 & 12/32 & 48/64 \\
 \hdashline
 5/2 & 4/16 & 20/32 & 100/64 \\
 \hdashline
 6/2 & 3/16 & 18/32 & 108/64 \\
 \hdashline
 7/2 & 2/16 & 14/32 & 98/64 \\
 \hdashline
 8/2 & 1/16 & 8/32 & 64/64 \\
 \hdashline
\end{array}"



D.

Mean of population 

"\mu=\dfrac{1+2+3+4}{4}=2.5"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=2.5=\mu"


Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{4}(2.25+0.25+0.25"


"+2.25)=1.25"

The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{440}{64}-(2.5)^2=0.625= \dfrac{\sigma^2}{n}"